# The $C^{0,1}$ Itô-Ventzell formula for weak Dirichlet processes

### v2

@@ Abstract
This paper proves an extension of the Itô-Ventzell formula that applies to stochastic flows in $C^{0,1}$ for continuous weak Dirichlet processes.
@@ 1 Introduction
where $W$ is a Brownian Motion. Moreover, we assume that $F$ is a continuous progressively measurable process in $t$ for fixed $x$ and twice differentiable in $x$ for fixed $t$ and $\omega$. Note that the dependence of $F$, $\beta$, and $\gamma$ on $\omega$ is suppressed. Then the Itô-Ventzell formula states that $F(t,X_{t})$ is also a continuous semimartingale and has the following decomposition:
@@
The Itô-Ventzell formula links stochastic processes to stochastic partial differential equations (SPDEs) and has various applications. It is used, for instance, to analyze various optimal control problems, to determine localization errors of SPDEs, for generalized stochastic forward integrals, for forward performance processes, and to develop pricing models.
@@ 2 Basic Definitions
Our processes may not be semimartingales and may not admit a quadratic (co-)variation. For this purpose, we need generalizations of these concepts, see Gozzi and Russo [12, p.1567]:
@@
Definition 2.2 ensures that the covariation and the forward integral are again continuous processes and that the covariation is bilinear and symmetric. In particular, the (forward) integral against the covariation is well-defined. For a more detailed explanation of the covariation and the integral against it, see Russo and Vallois [21, pp. 83-97].
@@
Let $X$ be a continuous process. If $[X]_{t}<\infty$ for all $t\in[0,T]$, then we call $X$ a finite quadratic variation (FQV) process.
The following definition is standard, see, for instance, Russo and Vallois [19, 20, 21, 21, 22, 23], Coquet et al. [5], Bandini and Russo [1, 2] or Gozzi and Russo [12, p.1570]:
@@
A real continuous process $X$ is called a weak Dirichlet process if $X$ is decomposable into a local martingale $M$ and a weak zero energy process $A$, i.e., $X=M+A$.
@@
The following processes are weak-Dirichlet processes according to Gozzi and Russo [12, pp.1570-1571], and Errami and Russo [9, p.267]:
@@
Contrary to the classical Itô formula, the Itô-Ventzell formula considers stochastic flows, which are defined in the following definition. Similar, but slightly different definitions can be found, e.g., in Kohlmann et al. [14, p.7] or dos Reis and Platonov [7, p.6].
@@
    $\beta:\Omega\times\mathbb{R}_{\geq 0}\times\mathbb{R}^{d}\to\mathbb{R}$ and $\gamma:\Omega\times\mathbb{R}_{\geq 0}\times\mathbb{R}^{d}\to\mathbb{R}^{d}$ are progressively measurable and previsible random fields,
@@ 3 Main Result
The assumptions of Theorem 3.1 are satisfied (through Assumption 1a) if the mappings $(t,x)\mapsto\beta(t,x)$ and $(t,x)\mapsto\gamma(t,x)$ are a.s. continuous, $\sup_{t\in[0,T]}|\gamma(t,0)|$ is $p$-integrable for all $p\in[1,\infty)$ and $x\mapsto\gamma(\cdot,x)$ is a.s. locally Lipschitz continuous with a Lipschitz constant independent of $t$ which is integrable in $\omega$.
The theorem generalizes the result of Gozzi and Russo [12, pp.1571-1574] to stochastic flows.
@@
By recalling Proposition 1.2 of Russo and Vallois [21, p.89] it holds:
@@
Let $X$ and $Y$ be two continuous FQV processes and $H$ be a càdlàg process defined on the interval $[0,T]$. Then, it holds:
@@
Let $X$ be a FQV process, $t\in[0,T]$, and $Z_{r}^{\varepsilon}$ be a progressively measurable process for all $\varepsilon>0$ such that $\int_{0}^{T}\left|Z_{r}^{\varepsilon}\right|^{2}\mathrm{d}r\xrightarrow[\mathbb{P}]{\varepsilon\to 0}0$. Then:
Proof. We get with Cauchy-Schwarz:
$\sup_{t\in[0,T]}\left|\int_{0}^{t}Z^{\varepsilon}_{r}\frac{X_{r+\varepsilon}-X_{r}}{\sqrt{\varepsilon}}\mathrm{d}r\right|\leq\int_{0}^{T}\left|Z^{\varepsilon}_{r}\frac{X_{r+\varepsilon}-X_{r}}{\sqrt{\varepsilon}}\right|\mathrm{d}r\leq\sqrt{\int_{0}^{T}\left|Z^{\varepsilon}_{r}\right|^{2}\mathrm{d}r}\sqrt{\int_{0}^{T}\frac{(X_{r+\varepsilon}-X_{r})^{2}}{\varepsilon}\mathrm{d}r}.$
@@
Let $X$ be a continuous FQV process, $t\in[0,T]$, and $Z_{r}^{\varepsilon}$ be a progressively measurable process for all $\varepsilon>0$ such that $\int_{0}^{T}|Z_{r}^{\varepsilon}|\mathrm{d}r$ is bounded in probability for $\varepsilon>0$. Then:
@@
Let $X$, $Y$ be two FQV processes, $t\in[0,T]$. Moreover, let $Z_{r}^{\varepsilon}$ be a progressively measurable process for all $\varepsilon>0$ such that $\sup_{r\in[0,T]}\left|Z_{r}^{\varepsilon}\right|\xrightarrow[\mathbb{P}]{\varepsilon\to 0}0$. Then:
Proof. We get:
and the lemma follows from Lemma 4.1.
@@
    If $F(t,x)\in C^{0,2}$ and $\gamma(t,x),\beta(t,x)\in C^{0,1}$ with $\alpha$-Hölder-continuous derivatives in $x$, then Flandoli and Russo [10, p.279] give us the Itô-Ventzell formula for weak Dirichlet processes:
@@
    Next note that $\int_{0}^{t}F_{x}(r,X_{r})\mathrm{d}^{-}X_{r}$ exists. Due to $\int_{0}^{t}F_{x}(r,X_{r})\mathrm{d}^{-}M_{r}=\int_{0}^{t}F_{x}(r,X_{r})\mathrm{d}M_{r}$ coinciding with the classical stochastic integral, $\int_{0}^{t}F_{x}(r,X_{r})\mathrm{d}^{-}A_{r}$ also exists. Moreover, when carefully scanning the to our setting adapted proof of the Itô-Ventzell formula in Flandoli and Russo [10, pp.279-280], we can reduce the regularity assumptions of $\beta$ and $\gamma$ due to Lemma 4.6. Thus, it is sufficient to assume that $\beta(t,x)\in C^{0,1}$ (without Hölder-continuity of $\beta_{x}(t,x)$ in $x$) and $\gamma(t,x)\in C^{0,1+\alpha}$ (with only local $\alpha$-Hölder-continuity of $\gamma_{x}(t,x)$ in $x$) in case (b).
@@
    First of all, since $X$ is continuous, we can use a classical localization argument to choose a compact set $K\subset\mathbb{R}$ such that $X_{t}(\omega)\in K$ for all $t\in[0,T]$ and almost all $\omega\in\Omega$. Then, by using again classical localization arguments, we can assume that $\sup_{x\in K}|\beta(t,x)|\in\mathcal{L}^{1}(\Omega\times[0,T],\mathrm{d}\mathbb{P}\times\mathrm{d}t)$, $\sup_{x\in K}|\gamma(t,x)|\in\mathcal{L}^{2}(\Omega\times[0,T],\mathrm{d}\mathbb{P}\times\mathrm{d}t)$ both from Definition 2.7(e). Now, the prerequisites of Lemma 4.6 are fulfilled, and, in combination with the Assumption 1b and Lemma 4.7, all of the (stochastic) integrals in this proof are well-defined.
@@
    Let $\omega\in\Omega$ be arbitrary such that $r\to X_{r}(\omega)$ is continuous. Hence, there exists a convex compact set $K$ such that $X_{r}(\omega)\in K$ for all $r\in[0,T]$. Since $F_{x}(t,x)$ is continuous, we know that $F_{x}(t,x)$ is uniformly continuous on $[0,T]\times K$. Thus, for all $L>0$ there exists a $0<\rho_{\omega}\leq 1$ such that it holds for all $r\in[0,T]$ and for all $0<\varepsilon\leq\rho_{\omega}$: $\left|Y_{r}^{\varepsilon}(\omega)\right|\leq L$. (Note that $\left\lVert(r+\varepsilon,X_{r}(\omega))-(r,X_{r}(\omega))\right\rVert_{2}=\varepsilon\leq\rho_{\omega}$.) In particular, we get for any arbitrary $L>0$ a $0<\rho_{\omega}\leq 1$ such that $\sup_{r\in[0,T]}\left|Y_{r}^{\varepsilon}(\omega)\right|\leq L$ for all $0<\varepsilon\leq\rho_{\omega}$. Hence, $\sup_{r\in[0,T]}\left|Y_{r}^{\varepsilon}(\omega)\right|\xrightarrow[a.s.]{\varepsilon\to 0}0$ and therefore also in probability.
    Next, note that it holds that $X_{r}(\omega)+\lambda(X_{r+\varepsilon}(\omega)-X_{r}(\omega))\in K$ for all $\lambda\in[0,1]$ since $X_{r}(\omega),X_{r+\varepsilon}(\omega)\in K$. Moreover, $X_{r}(\omega)$ is uniformly continuous in $r$ for $r\in[0,T]$, i.e., for all $\tilde{L}>0$ there exists a $0<\tilde{\rho}_{\omega}\leq 1$ such that for all $r\in[0,T]$ and for all $0<\varepsilon\leq\tilde{\rho}_{\omega}$ it holds: $\left|X_{r+\varepsilon}(\omega)-X_{r}(\omega)\right|\leq\tilde{L}$. Let $L>0$ arbitrary. Since $F_{x}(t,x)$ is uniformly continuous on $[0,T]\times K$, there exists a $\delta>0$ such that for all $x,y\in K$ with $\left\lVert(t,x)-(t,y)\right\rVert_{2}\leq\delta$ it holds that $\left|F_{x}(t,x)-F_{x}(t,y)\right|\leq L$ for all $t\in[0,T]$ fixed. Now choose $\tilde{L}=\delta$ and define $\rho_{\omega}:=\min\{\delta,\tilde{\rho}_{\omega}\}$. Then
@@
    for all $0<\varepsilon\leq\rho_{\omega}$, all $\lambda\in[0,1]$, and all $r\in[0,T]$. This implies that $\left|Z_{r}^{\varepsilon}(\omega)\right|\leq L$ for all $0<\varepsilon\leq\rho_{\omega}$ and all $r\in[0,T]$. In particular, we get $\sup_{r\in[0,T]}\left|Z_{r}^{\varepsilon}(\omega)\right|\leq L$ for all $0<\varepsilon\leq\rho_{\omega}$. Since we chose $L>0$ and $\omega\in\Omega$ arbitrary, $\sup_{r\in[0,T]}\left|Z_{r}^{\varepsilon}\right|\xrightarrow[a.s.]{\varepsilon\to 0}0$ and therefore also in probability. Now, Lemma 4.5 yields that $\lim_{\varepsilon\to 0}R_{1,1}(t,\varepsilon)=0$ and $\lim_{\varepsilon\to 0}R_{1,2}(t,\varepsilon)=0$ in $ucp$.
@@
    In the next step, we want to show that this convergence also holds in $L^{2}(\mathrm{d}\mathbb{P})$. For this purpose, we show that $|\widetilde{Z}^{\varepsilon}_{r}|^{2}$ is uniformly integrable. Let $\varrho>0$ be arbitrary. Then, we get using Mao[17, p.39] (with $p=2+\varrho$):
    $\mathbb{E}|\widetilde{Z}^{\varepsilon}_{r}|^{2+\varrho}$ $=\displaystyle\frac{1}{\varepsilon^{1+\frac{\varrho}{2}}}\mathbb{E}\left|\int_{r}^{r+\varepsilon}\left[\gamma(s,X_{r})-\gamma(r,X_{r})\right]\mathrm{d}W_{s}\right|^{2+\varrho}$
@@
    $\mathbb{E}|\widetilde{Z}^{\varepsilon}_{r}|^{2+\varrho}$ $\leq\tilde{C}\cdot\displaystyle\frac{1}{\varepsilon}\int_{r}^{r+\varepsilon}\mathbb{E}\left|\gamma(s,X_{r})-\gamma(r,X_{r})\right|^{2+\varrho}\mathrm{d}s$
    $\leq\tilde{C}\sup_{s\in[0,T]}\mathbb{E}\left|\gamma(s,X_{r})-\gamma(r,X_{r})\right|^{2+\varrho}\leq C\sup_{s\in[0,T]}\mathbb{E}\left|\gamma(s,X_{r})\right|^{2+\varrho}\leq C\cdot B_{2}.$ (4.1)
    Thus, $\widetilde{Z}^{\varepsilon}_{r}\xrightarrow[\mathcal{L}^{2}(\mathrm{d}\mathbb{P})]{\varepsilon\to 0}0$ for all $r\in[0,T]$. Analogously to (Proof of Theorem 3.1.), we also get $\sup_{r\in[0,T]}\mathbb{E}|\widetilde{Z}^{\varepsilon}_{r}|^{2}\leq C\cdot B_{2}$. Hence, the bounded convergence theorem gives us that $\int_{0}^{T}\mathbb{E}|\widetilde{Z}^{\varepsilon}_{r}|^{2}\mathrm{d}r\xrightarrow{\varepsilon\to 0}0$. Thus, $\int_{0}^{T}|\widetilde{Z}^{\varepsilon}_{r}|^{2}\mathrm{d}r\xrightarrow[\mathbb{P}]{\varepsilon\to 0}0$ and Lemma 4.2 implies that $\lim_{\varepsilon\to 0}R_{2,2}=0$ in $ucp$ which gives us the proof of the Theorem. ∎

### v3

@@ Abstract
This paper proves an extension of the Itô-Ventzell formula that applies to stochastic flows in $C^{0,1}$ for continuous weak Dirichlet processes. We apply this theorem, for example, to give a representation result for strong solutions of time-dependent elliptic SPDEs, to derive formulas for quadratic variations, and to relax assumptions in a financial mathematics context.
@@ 1 Introduction
where $W$ is a Brownian Motion. Moreover, we assume that $F$ is a continuous progressively measurable process for fixed $x$ and twice differentiable in $x$ for fixed $t$ and $\omega$. Note that the dependence of $F$, $\beta$, and $\gamma$ on $\omega$ is suppressed. Then the Itô-Ventzell formula states that $F(t,X_{t})$ is also a continuous semimartingale and has the following decomposition:
@@
The Itô-Ventzell formula links stochastic processes to stochastic partial differential equations (SPDEs) and has various applications. It is used, for instance, to analyze various optimal control problems, to determine localization errors of SPDEs, for generalized stochastic forward integrals, for forward performance processes, and to develop pricing models. We apply our new version of the Itô-Ventzell formula to derive stochastic calculus formulas and to obtain a representation result for strong solutions of time-dependent stochastic elliptic PDEs. Here, we generalize naturally the solution theory of SPDEs to the setting of the paper. Moreover, we apply our result in a financial mathematics context to obtain the dynamics of the portfolio of a large investor whose trading influences prices.
@@ 2 Basic Definitions
Our processes may not be semimartingales and may not admit a quadratic (co-)variation. For this purpose, we need generalizations of these concepts, see Gozzi and Russo [13, p.1567]:
@@
Definition 2.2 ensures that the covariation and the forward integral are again continuous processes and that the covariation is bilinear and symmetric. In particular, the (forward) integral against the covariation is well-defined. For a more detailed explanation of the covariation and the integral against it, see Russo and Vallois [22, pp. 83-97].
@@
Let $X$ be a continuous and adapted process. If $[X]_{t}<\infty$ for all $t\in[0,T]$, then we call $X$ a finite quadratic variation (FQV) process.
The following definition is standard, see, for instance, Russo and Vallois [20, 21, 22, 22, 23, 24], Coquet et al. [6], Bandini and Russo [1, 2] or Gozzi and Russo [13, p.1570]:
@@
A real continuous and adapted process $X$ is called a weak Dirichlet process if $X$ is decomposable into a local martingale $M$ and a weak zero energy process $A$, i.e., $X=M+A$.
@@
The following processes are weak-Dirichlet processes according to Gozzi and Russo [13, pp.1570-1571], and Errami and Russo [10, p.267]:
@@
Contrary to the classical Itô formula, the Itô-Ventzell formula considers stochastic flows, which are defined in the following definition. Similar, but slightly different definitions can be found, e.g., in Kohlmann et al. [15, p.7] or dos Reis and Platonov [8, p.6].
@@
    $\beta:\Omega\times\mathbb{R}_{\geq 0}\times\mathbb{R}^{d}\to\mathbb{R}$ and $\gamma:\Omega\times\mathbb{R}_{\geq 0}\times\mathbb{R}^{d}\to\mathbb{R}^{d}$ are previsible random fields,
@@ 3 Main Result
In the last two integrals of Theorem 3.1 (b) both the integrand and the integrator are matrices. These integrals are, as usual, defined as the trace of the matrix of the corresponding integrals. The assumptions of Theorem 3.1 are satisfied (through Assumption 1a) if the mappings $(t,x)\mapsto\beta(t,x)$ and $(t,x)\mapsto\gamma(t,x)$ are a.s. continuous, $\sup_{t\in[0,T]}|\gamma(t,0)|$ is $p$-integrable for all $p\in[1,\infty)$ and $x\mapsto\gamma(\cdot,x)$ is a.s. locally Lipschitz continuous with a Lipschitz constant independent of $t$ which is integrable in $\omega$.
The theorem generalizes the result of Gozzi and Russo [13, pp.1571-1574] to stochastic flows.
@@
By recalling Proposition 1.2 of Russo and Vallois [22, p.89] it holds:
@@
Let $X$ and $Y$ be two FQV processes and $H$ be a càdlàg process defined on the interval $[0,T]$. Then, it holds:
@@
Let $X$, $Y$ be two FQV processes, $t\in[0,T]$. Moreover, let $Z^{\varepsilon}$ be a measurable process in $r$ for all $\varepsilon>0$ such that $\sup_{r\in[0,T]}\left|Z_{r}^{\varepsilon}\right|\xrightarrow[a.s.]{\varepsilon\to 0}0$. Then:
Proof. First, we note that $Z^{\varepsilon}$ is a.s. bounded for all $\varepsilon>0$ (with a bound possibly depending on $\omega$), i.e., the integral is well-defined. Then, we get:
and the lemma follows from Lemma 4.1.
@@
Let $X$ be a continuous FQV process, $t\in[0,T]$, and $Z^{\varepsilon}$ be a measurable and a.s. bounded process for all $\varepsilon>0$ such that $\int_{0}^{T}|Z_{r}^{\varepsilon}|\mathrm{d}r$ is bounded in probability in $\varepsilon>0$. Then:
@@
Let $X$ be a FQV process, $t\in[0,T]$, and $Z^{\varepsilon}$ be a measurable and a.s. bounded process for all $\varepsilon>0$ such that $\int_{0}^{T}\left|Z_{r}^{\varepsilon}\right|^{2}\mathrm{d}r\xrightarrow[\mathbb{P}]{\varepsilon\to 0}0$. Then:
Proof. We get with Cauchy-Schwarz:
$\sup_{t\in[0,T]}\left|\int_{0}^{t}Z^{\varepsilon}_{r}\frac{X_{r+\varepsilon}-X_{r}}{\sqrt{\varepsilon}}\mathrm{d}r\right|\leq\int_{0}^{T}\left|Z^{\varepsilon}_{r}\frac{X_{r+\varepsilon}-X_{r}}{\sqrt{\varepsilon}}\right|\mathrm{d}r\leq\sqrt{\int_{0}^{T}\left|Z_{r}^{\varepsilon}\right|^{2}\mathrm{d}r}\sqrt{\int_{0}^{T}\frac{(X_{r+\varepsilon}-X_{r})^{2}}{\varepsilon}\mathrm{d}r}.$
@@
    If $F(t,x)\in C^{0,2}$ and $\gamma(t,x),\beta(t,x)\in C^{0,1}$ with $\alpha$-Hölder-continuous derivatives in $x$, then Flandoli and Russo [11, p.279] give us the Itô-Ventzell formula for weak Dirichlet processes:
@@
    Next note that $\int_{0}^{t}F_{x}(r,X_{r})\mathrm{d}^{-}X_{r}$ exists. Due to $\int_{0}^{t}F_{x}(r,X_{r})\mathrm{d}^{-}M_{r}=\int_{0}^{t}F_{x}(r,X_{r})\mathrm{d}M_{r}$ coinciding with the classical stochastic integral, $\int_{0}^{t}F_{x}(r,X_{r})\mathrm{d}^{-}A_{r}$ also exists. Moreover, when carefully scanning the to our setting adapted proof of the Itô-Ventzell formula in Flandoli and Russo [11, pp.279-280], we can reduce the regularity assumptions of $\beta$ and $\gamma$ due to Lemma 4.6. Thus, it is sufficient to assume that $\beta(t,x)\in C^{0,1}$ (without Hölder-continuity of $\beta_{x}(t,x)$ in $x$) and $\gamma(t,x)\in C^{0,1+\alpha}$ (with only local $\alpha$-Hölder-continuity of $\gamma_{x}(t,x)$ in $x$) in case (b).
@@
    First of all, since $X$ is continuous, we can use a classical localization argument and assume that there exists a compact set $K\subset\mathbb{R}$ such that $X_{t}(\omega)\in K$ for all $t\in[0,T]$ and almost all $\omega\in\Omega$. Then, by using again classical localization arguments, we can assume that $\sup_{x\in K}|\beta(t,x)|\in\mathcal{L}^{1}(\Omega\times[0,T],\mathrm{d}\mathbb{P}\times\mathrm{d}t)$, $\sup_{x\in K}|\gamma(t,x)|\in\mathcal{L}^{2}(\Omega\times[0,T],\mathrm{d}\mathbb{P}\times\mathrm{d}t)$ both from Definition 2.7(e). Now, the prerequisites of Lemma 4.6 are fulfilled, and thus all of the (stochastic) integrals in this proof are well-defined.
@@
    Let $\omega\in\Omega$ be arbitrary such that $r\to X_{r}(\omega)$ is continuous. Hence, there exists a convex compact set $K$ such that $X_{r}(\omega)\in K$ for all $r\in[0,T]$. Since $F_{x}(t,x)$ is continuous, we know that $F_{x}(t,x)$ is uniformly continuous on $[0,T]\times K$. Thus, for all $L>0$ there exists a $0<\rho_{\omega}\leq 1$ such that it holds for all $r\in[0,T]$ and for all $0<\varepsilon\leq\rho_{\omega}$: $\left|Y_{r}^{\varepsilon}(\omega)\right|\leq L$. (Note that $\left\lVert(r+\varepsilon,X_{r}(\omega))-(r,X_{r}(\omega))\right\rVert_{2}=\varepsilon\leq\rho_{\omega}$.) In particular, we get for any arbitrary $L>0$ a $0<\rho_{\omega}\leq 1$ such that $\sup_{r\in[0,T]}\left|Y_{r}^{\varepsilon}(\omega)\right|\leq L$ for all $0<\varepsilon\leq\rho_{\omega}$. Hence, $\sup_{r\in[0,T]}\left|Y_{r}^{\varepsilon}(\omega)\right|\xrightarrow[a.s.]{\varepsilon\to 0}0$.
    Next, note that it holds that $X_{r}(\omega)+\lambda(X_{r+\varepsilon}(\omega)-X_{r}(\omega))\in K$ for all $\lambda\in[0,1]$ since $X_{r}(\omega),\linebreak X_{r+\varepsilon}(\omega)\in K$. Moreover, $X_{r}(\omega)$ is uniformly continuous in $r$ for $r\in[0,T]$, i.e., for all $\tilde{L}>0$ there exists a $0<\tilde{\rho}_{\omega}\leq 1$ such that for all $r\in[0,T]$ and for all $0<\varepsilon\leq\tilde{\rho}_{\omega}$ it holds: $\left|X_{r+\varepsilon}(\omega)-X_{r}(\omega)\right|\leq\tilde{L}$. Let $L>0$ arbitrary. Since $F_{x}(t,x)$ is uniformly continuous on $[0,T]\times K$, there exists a $\delta>0$ such that for all $x,y\in K$ with $\left\lVert(t,x)-(t,y)\right\rVert_{2}\leq\delta$ it holds that $\left|F_{x}(t,x)-F_{x}(t,y)\right|\leq L$ for all $t\in[0,T]$ fixed. Now choose $\tilde{L}=\delta$ and define $\rho_{\omega}:=\min\{\delta,\tilde{\rho}_{\omega}\}$. Then
@@
    for all $0<\varepsilon\leq\rho_{\omega}$, all $\lambda\in[0,1]$, and all $r\in[0,T]$. This implies that $\left|Z_{r}^{\varepsilon}(\omega)\right|\leq L$ for all $0<\varepsilon\leq\rho_{\omega}$ and all $r\in[0,T]$. In particular, we get $\sup_{r\in[0,T]}\left|Z_{r}^{\varepsilon}(\omega)\right|\leq L$ for all $0<\varepsilon\leq\rho_{\omega}$. Since we chose $L>0$ and $\omega\in\Omega$ arbitrary, $\sup_{r\in[0,T]}\left|Z_{r}^{\varepsilon}\right|\xrightarrow[a.s.]{\varepsilon\to 0}0$. Now, Lemma 4.2 yields that $\lim_{\varepsilon\to 0}R_{1,1}(t,\varepsilon)=0$ and $\lim_{\varepsilon\to 0}R_{1,2}(t,\varepsilon)=0$ in $ucp$.
@@
    First, we note that both $\widetilde{Y}^{\varepsilon}$ and $\widetilde{Z}^{\varepsilon}$ are a.s. bounded for all $\varepsilon>0$ since $\sup_{x\in K}|\beta(t,x)|\in\mathcal{L}^{1}(\Omega\times[0,T],\mathrm{d}\mathbb{P}\times\mathrm{d}t)$, $\sup_{x\in K}|\gamma(t,x)|\in\mathcal{L}^{2}(\Omega\times[0,T],\mathrm{d}\mathbb{P}\times\mathrm{d}t)$, and $X$ is a.s. continuous.
@@
    In the next step, we want to show that this convergence also holds in $L^{2}(\mathrm{d}\mathbb{P})$. For this purpose, we show that $|\widetilde{Z}^{\varepsilon}_{r}|^{2}$ is uniformly integrable. Let $\varrho>0$ be arbitrary. Then, we get using Mao[18, p.39] (with $p=2+\varrho$):
    $\mathbb{E}[|\widetilde{Z}^{\varepsilon}_{r}|^{2+\varrho}]$ $=\displaystyle\frac{1}{\varepsilon^{1+\frac{\varrho}{2}}}\mathbb{E}[\left|\int_{r}^{r+\varepsilon}\left[\gamma(s,X_{r})-\gamma(r,X_{r})\right]\mathrm{d}W_{s}\right|^{2+\varrho}]$
@@
    $\mathbb{E}[|\widetilde{Z}^{\varepsilon}_{r}|^{2+\varrho}]$ $\leq\tilde{C}\cdot\displaystyle\frac{1}{\varepsilon}\int_{r}^{r+\varepsilon}\mathbb{E}\left|\gamma(s,X_{r})-\gamma(r,X_{r})\right|^{2+\varrho}\mathrm{d}s$
    $\leq\tilde{C}\sup_{s\in[0,T]}\mathbb{E}[\left|\gamma(s,X_{r})-\gamma(r,X_{r})\right|^{2+\varrho}]\leq C\sup_{s\in[0,T]}\mathbb{E}[\left|\gamma(s,X_{r})\right|^{2+\varrho}]\leq C\cdot B_{2}.$ (4.1)
    Thus, $\widetilde{Z}^{\varepsilon}_{r}\xrightarrow[\mathcal{L}^{2}(\mathrm{d}\mathbb{P})]{\varepsilon\to 0}0$ for all $r\in[0,T]$. Analogously to (Proof of Theorem 3.1.), we also get $\sup_{r\in[0,T]}\mathbb{E}[|\widetilde{Z}^{\varepsilon}_{r}|^{2}]\leq C\cdot B_{2}$. Hence, the bounded convergence theorem gives us that $\int_{0}^{T}\mathbb{E}|\widetilde{Z}^{\varepsilon}_{r}|^{2}\mathrm{d}r\xrightarrow{\varepsilon\to 0}0$. Thus, $\int_{0}^{T}|\widetilde{Z}^{\varepsilon}_{r}|^{2}\mathrm{d}r\xrightarrow[\mathbb{P}]{\varepsilon\to 0}0$ and Lemma 4.5 implies that $\lim_{\varepsilon\to 0}R_{2,2}=0$ in $ucp$ which gives us the proof of the Theorem. ∎